OpenStudy (anonymous):

find the sum of the infinite geometric series 3-sqrt(3)+1-

OpenStudy (anonymous):

\[3-\sqrt{3}+1-...\]

OpenStudy (anonymous):

\[ =3\cdot \sum^\infty_{k=0} \left(-\frac{1}{\sqrt{3}}\right)^k \] Use the formula for geometric series!

OpenStudy (anonymous):

\[ = \frac{3}{1+1/\sqrt{3}} = 9\, \left( 3+\sqrt {3} \right) ^{-1} \]